OpenStudy (anonymous):

25^-3/2 equals? and (1/8)^4/3

jhonyy9 (jhonyy9):

25=5^2 25^(-3/2) = 1/(5^2)^(3/2) = 1/5^(2*3/2) = 1/5^3 =1/125

jhonyy9 (jhonyy9):

8=2^3 (1/8)^(4/3) = 1/(2^3)^(4/3) =1/2^4 = 1/16

OpenStudy (anonymous):

thanks jhonyy9

jhonyy9 (jhonyy9):

yw good luck bye